# Cluster Magnification, Root Capacity, Unique Chains and Base Change

### v3

@@ 4.2. Strong cluster magnification problem for irreducible polynomials
    By Prop. 4.1.9, if for some $i$, $M_{i}/K$ is obtained by strong cluster magnification from $L_{i}/K$ then for each $1\leq j\leq s$, $M_{j}/K$ is obtained by strong cluster magnification from some subextension $L_{j}/K$.
@@
$g$ over $K$ is obtained by nontrivial cluster magnification from an $f$ over $K$ if and only if ${{\rm Gal}}(K_{g}/K)\cong A\times B$ for nontrivial subgroups $A$ and $B$ and (for $\gamma$ as above) ${{\rm Gal}}(K_{g}/K(\gamma))\cong A^{\prime}\times 1$ (under the same isomorphism) for a subgroup $A^{\prime}\subset A$ with $[A:A^{\prime}]>2$ and $f$ is the minimal polynomial for a primitive element of $(K_{g})^{A^{\prime}\times B}$ over $K$.
@@
An irreducible polynomial $g$ over $K$ is called primitive if it is not obtained by a nontrivial strong cluster magnification over $K$. (this notion occurs in [5] as well).
@@
Some simple examples / cases of primitive polynomials $g$ over $K$ follow from Thm. 2.2.3 and Thm. 4.2.3.
@@ 4.3. Weak cluster magnification
Let $L_{k}$ be as in proof of Thm. 3.2.4. Then for $j<k$, $L_{k}/K$ is obtained by nontrivial weak cluster magnification from $L_{j}/K$ but is not obtained by nontrivial strong cluster magnification from $L_{j}/K$ since $\tilde{L_{j}}\cap L_{k}=L_{k}\neq L_{j}$.
@@
    Let $n,l$ be integers such that $6<l<n$, $l|n$ such that $n=lm$ and $\text{gcd}(l,m)=1$. Let $M=\mathbb{Q}(\xi_{n})$ and $L=\mathbb{Q}(\xi_{l})$. We have
@@ 5.1. Cluster tower of a polynomial
By [3, Prop. 1 and Thm. A ], the Galois group $G$ of splitting field of $f$ is isomorphic to $\mathbb{Z}/n\mathbb{Z}\rtimes(\mathbb{Z}/n\mathbb{Z})^{\times}$ if and only if $n$ is odd or, $n$ is even with $\sqrt{c}\not\in\mathbb{Q}(b)$ if and only if $\mathbb{Q}(c^{1/n})\cap\mathbb{Q}(b)=\mathbb{Q}$. Assume $n$ to satisfy these conditions. Hence, in particular the order of $G$ is $n\phi(n)$.
@@
Because of our assumptions, we have $m+1\leq s$. Let the representatives of clusters of roots of $f$ be given by
$\beta_{3}=ab^{m},\beta_{m+1}=ab^{2},\beta_{i}=ab^{i-1}\ \text{ for other}\ 1\leq i\leq s.$
@@ 5.2. Group theoretic formulation of cluster towers
Since $\dfrac{a_{i}}{a_{i-1}}\leq\left(n-(m_{i}-1)r\right)$, for all $i\geq 1$ (where $r=r_{K}(f)$), we have
@@
Now since $H_{i}=\sigma_{i}H\sigma_{i}^{-1}$ for all $i$. Hence $N_{G}(H_{i})=\sigma_{i}N_{G}(H)\sigma_{i}^{-1}.$ From Thm. 2.1.1 (3), we have $r_{K}(K_{m})=[N_{G}(J_{m}):J_{m}]$.
@@ 6.1. The group of automorphisms of finite extensions
Proof of (3): This corresponds to the homomorphism $\Phi$ being surjective.
@@ 6.2. Root capacity
We also have $\alpha=\lambda\circ\mu(\alpha)$. Hence $f(x)|(\lambda\circ\mu(x)-x)$. Thus $\lambda\circ\mu(\alpha_{i})=\alpha_{i}$ for all $i$. Hence $\mu$ is a bijection from the set of roots of $f$ to set of roots of $g$. By relabelling we can assume $\mu(\alpha_{i})=\beta_{i}$ for all $i$. Thus $\lambda(\beta_{i})=\alpha_{i}$ for all $i$. Hence for any $i$, we have $\alpha_{i}\in M\iff\beta_{i}\in M$.
@@
Proof of (2): Let $L=K(\alpha)$ for $\alpha\in\bar{K}$ with $f$ as the minimal polynomial of $\alpha$ over $K$ with roots $\{\alpha_{i}\}_{1\leq i\leq n}$. Then $\rho_{K}(M,L)=r_{K}(L)$ is equivalent to the statement  $\alpha_{i}\in M\iff\alpha_{i}\in L$. If $\sigma\in{\rm Aut}(M/K)$, then $\sigma|_{L}$ maps $\alpha$ to an $\alpha_{i}\in M$. Hence, $\sigma|_{L}(\alpha)\in L$. Thus $\sigma|_{L}\in{\rm Aut}(L/K)$.
@@
Let $\alpha\in\bar{K}$. Then for any $M/K$ we have $r_{K}(f)|\rho_{K}(M,\alpha)$. Hence for any $L/K$, we have $r_{K}(L)|\rho_{K}(M,L)$. That is $\rho_{K}(M,L)=a.r_{K}(L)$ where $0\leq a\leq s_{K}(L)$.
@@
Let notation be as in subsection 5.2. Let $L=K_{1}$ and $m_{0}=1$. Then for $i\geq 0$ and $m_{i}\leq m<m_{i+1}$, we have $\rho(K_{m},L)=\rho(K_{m_{i}},L)=(m_{i+1}-1)r_{K}(L)$.
@@
Consider the set of $s$ many representatives of clusters of minimal polynomial of a primitive element of $L/K$. Consider the subsets of representatives contained in $M_{1}$ and $M_{2}$ and let $a$ and $b$ be their respective cardinalities. Hence $\rho_{K}(M_{1})=a.r_{K}(L)$ and $\rho_{K}(M_{2})=b.r_{K}(L)$. Let $i$ be cardinality of intersection of these two sets. Hence $\rho_{K}(M_{1}\cap M_{2},L)\geq i.r_{K}(L)$. Also since cardinality of union of these two sets is $a+b-i$. Hence $\rho_{K}(M_{1}M_{2},L)\geq(a+b-i).r_{K}(L)$. Hence we are done.
@@
$r_{K}(M)|(r_{L}(M)r_{K}(L))$. Since $r_{K}(M)=[M:L]r_{K}(L)$, we conclude that $r_{L}(M)=[M:L]$ and $M/L$ is Galois.
@@ 7.3. Interesting Examples
Let $L_{1}$ be Galois closure of $L/F_{1}$. Hence $L_{1}=\mathbb{Q}(a,b^{2})$. Also since $\mathbb{Q}(a)\cap\mathbb{Q}(b)=\mathbb{Q}$. Hence $L\cap\mathbb{Q}(b^{2})=\mathbb{Q}$. By Lemma 2.2.2, $L\cap F_{1}(b^{2})=F_{1}$. Hence ${\rm Gal}(L_{1}/F_{1})$ can be identified with $\mathbb{Z}/(n/2)\mathbb{Z}\rtimes(\mathbb{Z}/(n/2)\mathbb{Z})^{\times}$. If $n/2$ is even, we can repeat the process for $F_{1}$ and get $F_{2}$. The process will terminate at $F_{v_{2}(n)}$ as $n/v_{2}(n)$ is odd.
@@
Let $f$ be an irreducible polynomial over number field $K$ with given $n>2$ and $1<r<n$ with $r|n$ as in proof of Thm. 3.1.1 and $s=n/r$. Let $L/K$ be extension formed by adjoining a root of $f$ to $K$. Then the unique descending chain is $L\supsetneq N\supsetneq K$ and the unique ascending chain is $K\subsetneq F\subsetneq L$ and they both coincide.
@@
Thus for $r<n$ (that is $s>1$), one can compute and show that $N_{G}(H)=H^{G}=({\mathbb{Z}}/r{\mathbb{Z}})^{s}\times\{0\}$. Hence by Prop. 7.2.2 (5), we are done.
@@
Let $L_{k}$ be as in proof of Thm. 3.2.4.
@@
    2. (2)
      We give a proof for case (a) only. The proof in the other case follows similarly. Now for any $l<n-1$, we have $\alpha_{i_{l+1}}\not\in K(\alpha_{i_{1}},\alpha_{i_{2}},\dots,\alpha_{i_{l}})$ and we also have $\alpha_{i_{n}}\in K(\alpha_{i_{1}},\alpha_{i_{2}},\dots,\alpha_{i_{n-1}})$ for distinct $i_{m}\leq n$. Since $t_{1}=\alpha_{1}+\alpha_{2}+\dots+\alpha_{k}$. Hence $t_{1}$ has at most $\ {}^{n}C_{k}$ conjugates inside $\tilde{L_{k}}$ of the form $\alpha_{i_{1}}+\alpha_{i_{2}}+\dots+\alpha_{i_{k}}$ for distinct $i_{m}\leq n$.
@@
  • Proof of last assertion: We have that the smallest normal subgroup of ${\mathfrak{S}}_{n}$ containing ${\mathfrak{S}}_{n-k}$ is ${\mathfrak{S}}_{n}$ itself. Hence by Thm. 7.2.1 (5), we are done.
@@ 8. Base Change theorems
In this section, we study the properties of cluster magnification and other phenomena after a special type of base change. First we see a result about field extensions.
The following lemma is a special case of [1, Prop. 2.5].
@@
Throughout this section, let $L/K$ be a finite extension and let $K^{\prime}/K$ be a finite Galois extension such that $\tilde{L}$ and $K^{\prime}$ are linearly disjoint over $K$. We consider base change by such extensions $K^{\prime}$.
8.1. A base change theorem for strong and weak cluster magnification
@@
Suppose $M/K$ is obtained by nontrivial strong cluster magnification from an $L/K$ with magnification factor $d$. Let $\tilde{M}$ and $K^{\prime}$ be also linearly disjoint over $K$. Then $MK^{\prime}/K^{\prime}$ is obtained by nontrivial strong cluster magnification from $LK^{\prime}/K^{\prime}$ with the same magnification factor $d$.
@@
    Let $F^{\prime}:=FK^{\prime}$. Since, $F/K$ is Galois, $F^{\prime}/K^{\prime}$ is Galois. Also, $[F^{\prime}:K^{\prime}]=[F:K]=d>1$. Let $\tilde{L^{\prime}}$ be Galois closure of $L^{\prime}$ over $K^{\prime}$ in $\bar{K}$. Now, $N:=\tilde{L}K^{\prime}$ is Galois over $K^{\prime}$ and $L^{\prime}\subset N$. Hence, $\tilde{L^{\prime}}\subset N$. Since $F$ is linearly disjoint with $\tilde{L}$ over $K$, we have $\tilde{L}\cap F=K$. Also $\tilde{L}\cap K^{\prime}=K$ and $F\cap K^{\prime}=K$. Since $\tilde{L},F$ and $K^{\prime}$ are Galois over $K$. From Lem. 8.0.1 for $E_{1}=\tilde{L},E_{2}=F,E_{3}=K^{\prime}$, we have $\tilde{L}F\cap K^{\prime}=K\iff\tilde{L}K^{\prime}\cap FK^{\prime}=K^{\prime}$. So we have $N\cap F^{\prime}=K^{\prime}$. Thus, $\tilde{L^{\prime}}\cap F=K^{\prime}$. Hence $\tilde{L^{\prime}}$ and $F^{\prime}$ are linearly disjoint over $K^{\prime}$.
@@
Lem. 8.1.6 gives an alternate proof for Thm. 8.1.1 by using the criterion in Thm. 4.1.6.
We conclude the following result for the strong cluster magnification of polynomials from Thm. 8.1.1.
@@
Suppose $g$ over $K$ is obtained by nontrivial strong cluster magnification from an $f$ over $K$ with magnification factor $d$. Let $K^{\prime}/K$, contained in $\bar{K},$ be Galois and linearly disjoint with $K_{g}$ over $K$. Then $g$ over $K^{\prime}$ is obtained by nontrivial strong cluster magnification from $f$ over $K^{\prime}$ with same magnification factor $d$.
@@
Suppose $M/K$ is obtained by weak cluster magnification from an $L/K$ with magnification factor $d$. Let $K^{\prime}/K$, contained in $\bar{K},$ be Galois and let $\tilde{M}$ and $K^{\prime}$ be linearly disjoint over $K$. Then $MK^{\prime}/K^{\prime}$ is obtained by weak cluster magnification from $LK^{\prime}/K^{\prime}$ with the same magnification factor $d$.
@@
Follows from last assertion in Lemma 8.1.6.

### v4

@@ 4.2. Strong cluster magnification problem for irreducible polynomials
    By Remark 4.1.9, if for some $i$, $M_{i}/K$ is obtained by strong cluster magnification from $L_{i}/K$ then for each $1\leq j\leq s$, $M_{j}/K$ is obtained by strong cluster magnification from some subextension $L_{j}/K$.
@@
A polynomial $g$ over $K$ is obtained by nontrivial cluster magnification from an $f$ over $K$ if and only if ${{\rm Gal}}(K_{g}/K)\cong A\times B$ for nontrivial groups $A$ and $B$ and (for $\gamma$ as above) ${{\rm Gal}}(K_{g}/K(\gamma))\cong A^{\prime}\times 1$ (under the same isomorphism) for a subgroup $A^{\prime}\subset A$ with $[A:A^{\prime}]>2$ and $f$ is the minimal polynomial for a primitive element of $(K_{g})^{A^{\prime}\times B}$ over $K$.
@@
An irreducible polynomial $g$ over $K$ is called polynomial with primitive clusters if it is not obtained by a nontrivial strong cluster magnification over $K$. (this notion occurs in [5] as well).
@@
Some simple examples / cases of polynomials with primitive clusters $g$ over $K$ follow from Thm. 2.2.3 and Thm. 4.2.3.
@@ 4.3. Weak cluster magnification
Let $f$ over $K$ be irreducible of deg $n$ with Galois group ${\mathfrak{S}}_{n}$ with roots $\alpha_{i}\in\bar{K}$ for $1\leq i\leq n$. For $21\leq k\leq n-$ 2, let $L_{k}=K(\alpha_{1},\dots,\alpha_{k})$. Then for $j<k$, $L_{k}/K$ is obtained by nontrivial weak cluster magnification from $L_{j}/K$ but is not obtained by nontrivial strong cluster magnification from $L_{j}/K$ since $\tilde{L_{j}}\cap L_{k}=L_{k}\neq L_{j}$.
@@
    Let $n,l$ be integers such that $6<l<n$, $l|n$ such that $n=lm$ and $\text{gcd}(l,m)=1$. Let $M=\mathbb{Q}(\xi_{n})$ and $L=\mathbb{Q}(\xi_{l})$. Since gcd$(l,m)=1$, we have $v_{p}(m)=v_{p}(n)$ for $p|m$. We have
@@ 5.1. Cluster tower of a polynomial
By [4, Prop. 1 and Thm. A ], the Galois group $G$ of splitting field of $f$ is isomorphic to $\mathbb{Z}/n\mathbb{Z}\rtimes(\mathbb{Z}/n\mathbb{Z})^{\times}$ if and only if $n$ is odd or, $n$ is even with $\sqrt{c}\not\in\mathbb{Q}(b)$ if and only if $\mathbb{Q}(c^{1/n})\cap\mathbb{Q}(b)=\mathbb{Q}$. Assume $n$ to satisfy these conditions. Hence, in particular the order of $G$ is $n\phi(n)$.
@@
Because of our assumptions, we have $m+1\leq s$. Let the representatives $\beta_{i}$s of clusters of roots of $f$, for $1\leq i\leq s$ be given by
$\beta_{3}=ab^{m},\ \beta_{m+1}=ab^{2},\ \text{and}\ \beta_{i}=ab^{i-1}\ \text{for}\ i\neq 3,m+1\ .$
@@ 5.2. Group theoretic formulation of cluster towers
Let $m_{0}=1$. Now for $i\geq 1$, $\dfrac{a_{i}}{a_{i-1}}=[K_{m_{i}}:K_{m_{i-1}}]$ and $K_{m_{i-1}}=K_{m_{i}-1}$ and number of roots of $f$ contained in $K_{m_{i}-1}$ is $\geq(m_{i}-1)r$ (where $r=r_{K}(f)$). Also $K_{m_{i}}=K_{m_{i}-1}(\beta_{m_{i}})$.
Thus $\dfrac{a_{i}}{a_{i-1}}=[K_{m_{i}}:K_{m_{i}-1}]\leq\left(n-(m_{i}-1)r\right)$ for all $i\geq 1$. Hence we have
@@
Now since $H_{i}=\sigma_{i}H\sigma_{i}^{-1}$ for all $i$, we have that $N_{G}(H_{i})=\sigma_{i}N_{G}(H)\sigma_{i}^{-1}.$ From Thm. 2.1.1 (3), we have $r_{K}(K_{m})=[N_{G}(J_{m}):J_{m}]$.
@@ 6.1. The group of automorphisms of finite extensions
Proof of (3): This corresponds to the homomorphism $\Phi$ being surjective. ∎
@@ 6.2. Root capacity
We also have $\alpha=\lambda\circ\mu(\alpha)$. Hence $f(x)|(\lambda\circ\mu(x)-x)$. Thus $\lambda\circ\mu(\alpha_{i})=\alpha_{i}$ for all $i$. Hence $\mu$ is a bijection from the set of roots of $f$ to set of roots of $g$. By relabelling we can assume $\mu(\alpha_{i})=\beta_{i}$ for all $i$. Thus $\lambda(\beta_{i})=\alpha_{i}$ for all $i$. Hence for any $i$, we have $\alpha_{i}\in M\iff\beta_{i}\in M$. ∎
@@
Proof of (2): Let $L=K(\alpha)$ for $\alpha\in\bar{K}$ with $f$ as the minimal polynomial of $\alpha$ over $K$ with roots $\{\alpha_{i}\}_{1\leq i\leq n}$. Then $\rho_{K}(M,L)=r_{K}(L)$ is equivalent to the statement  $\alpha_{i}\in M\iff\alpha_{i}\in L$. If $\sigma\in{\rm Aut}(M/K)$, then $\sigma|_{L}$ maps $\alpha$ to an $\alpha_{i}\in M$. Hence, $\sigma|_{L}(\alpha)\in L$. Thus $\sigma|_{L}\in{\rm Aut}(L/K)$. ∎
@@
Let $\alpha\in\bar{K}$ and let $f$ be minimal polynomial of $\alpha$ over $K$. Then for any $M/K$ we have $r_{K}(f)|\rho_{K}(M,\alpha)$. Hence for any $L/K$, we have $r_{K}(L)|\rho_{K}(M,L)$. That is $\rho_{K}(M,L)=a.r_{K}(L)$ where $0\leq a\leq s_{K}(L)$.
@@
Let notation be as in subsection 5.2. Let $L=K_{1}$ and $m_{0}=1$. Then for $i\geq 0$ and $m_{i}\leq m<m_{i+1}$, we have $\rho(K_{m},L)=\rho(K_{m_{i}},L)\geq(m_{i+1}-1)r_{K}(L)$.
@@
Consider the set of $s$ many representatives of clusters of minimal polynomial of a primitive element of $L/K$. Consider the subsets of representatives contained in $M_{1}$ and $M_{2}$ and let $a$ and $b$ be their respective cardinalities. Hence $\rho_{K}(M_{1})=a.r_{K}(L)$ and $\rho_{K}(M_{2})=b.r_{K}(L)$. Let $i$ be cardinality of intersection of these two sets. Hence $\rho_{K}(M_{1}\cap M_{2},L)\geq i.r_{K}(L)$. Also since cardinality of union of these two sets is $a+b-i$. Hence $\rho_{K}(M_{1}M_{2},L)\geq(a+b-i).r_{K}(L)$. Hence we are done. ∎
@@
$r_{K}(M)|(r_{L}(M)r_{K}(L))$. Since $r_{K}(M)=[M:L]r_{K}(L)$, we conclude that $r_{L}(M)=[M:L]$ and $M/L$ is Galois. ∎
@@ 7.3. Interesting Examples
Let $L_{1}$ be Galois closure of $L/F_{1}$. Hence $L_{1}=\mathbb{Q}(a,b^{2})$. Also since $\mathbb{Q}(a)\cap\mathbb{Q}(b)=\mathbb{Q}$. Hence $L\cap\mathbb{Q}(b^{2})=\mathbb{Q}$. By Lemma 2.2.2, $L\cap F_{1}(b^{2})=F_{1}$. Hence ${\rm Gal}(L_{1}/F_{1})$ can be identified with $\mathbb{Z}/(n/2)\mathbb{Z}\rtimes(\mathbb{Z}/(n/2)\mathbb{Z})^{\times}$. If $n/2$ is even, we can repeat the process for $F_{1}$ and get $F_{2}$. The process will terminate at $F_{v_{2}(n)}$ as $n/v_{2}(n)$ is odd. ∎
@@
Let $f$ be the irreducible polynomial over number field $K$ as in proof of Thm. 3.1.1 with given $n>2$ and $1<r<n$ with $r|n$ and $s=n/r$. Let $L/K$ be extension formed by adjoining a root of $f$ to $K$. Then the unique descending chain is $L\supsetneq N\supsetneq K$ and the unique ascending chain is $K\subsetneq F\subsetneq L$ and they both coincide.
@@
Thus for $r<n$ (that is $s>1$), one can compute and show that $N_{G}(H)=H^{G}=({\mathbb{Z}}/r{\mathbb{Z}})^{s}\times\{0\}$. Hence by Prop. 7.2.2 (5), we are done.∎
@@
Let $f$ over $K$ be irreducible of deg $n$ with Galois group ${\mathfrak{S}}_{n}$ with roots $\alpha_{i}\in\bar{K}$ for $1\leq i\leq n$. For $21\leq k\leq n-$ 2, let $L_{k}=K(\alpha_{1},\dots,\alpha_{k})$.
  1. (1)
@@
    2. (ii)
      We give a proof for case (a) only. The proof in the other case follows similarly. Now for any $l<n-1$, we have $\alpha_{i_{l+1}}\not\in K(\alpha_{i_{1}},\alpha_{i_{2}},\dots,\alpha_{i_{l}})$ and we also have $\alpha_{i_{n}}\in K(\alpha_{i_{1}},\alpha_{i_{2}},\dots,\alpha_{i_{n-1}})$ for distinct $i_{m}\leq n$. Since $t_{1}=\alpha_{1}+\alpha_{2}+\dots+\alpha_{k}$, we have that $t_{1}$ has at most $\ {}^{n}C_{k}$ conjugates inside $\tilde{L_{k}}$ of the form $\alpha_{i_{1}}+\alpha_{i_{2}}+\dots+\alpha_{i_{k}}$ for distinct $i_{m}\leq n$.
@@
  2. (2)
    Proof of last assertion: We have that the smallest normal subgroup of ${\mathfrak{S}}_{n}$ containing ${\mathfrak{S}}_{n-k}$ is ${\mathfrak{S}}_{n}$ itself. This is because $k\leq n-2$ that is $n-k\geq 2$. We know that for $n=3$ and $n\geq 5$, the only proper non trivial normal subgroup of ${\mathfrak{S}}_{n}$ is ${\mathfrak{A}}_{n}$ and ${\mathfrak{A}}_{n}$ cannot contain ${\mathfrak{S}}_{n-k}$ for any $k\leq n-2$. We also know that for $n=4$, the only proper non trivial normal subgroups of ${\mathfrak{S}}_{4}$ are ${\mathfrak{A}}_{4}$ and $V_{4}$ (Klein four-subgroup) and ${\mathfrak{A}}_{4}$ and $V_{4}$ cannot contain ${\mathfrak{S}}_{4-k}$ for both $k=1,2$. Hence by Thm. 7.2.1 (5), we are done.
@@ 8. Base Change theorems
All along we have been discussing clusters for irreducible polynomial over some fixed perfect base field $K$, or equivalently for extension $L/K$. Now we would like to define a notion of ‘base change’.
Definition 8.0.1.
We say that a finite extension $L/K$ has a ‘base change of $K$ by an extension $K^{\prime}/K$’ if $K^{\prime}/K$ is a finite Galois extension such that $\tilde{L}$ and $K^{\prime}$ are linearly disjoint over $K$.
Remark 8.0.2.
Consider $M/L/K$. If $M/K$ has a base change of $K$ by $K^{\prime}$ then $L/K$ also has a base change of $K$ by $K^{\prime}$.
Throughout this section we consider extensions $L/K$ which have base change of $K$ by $K^{\prime}$ and study the properties of cluster magnification and other phenomena for corresponding extensions $LK^{\prime}/K^{\prime}$.
The following result establishes a certain conditional transitivity of base change.
Proposition 8.0.3.
Consider an extension $L/K$ which has a base change of $K$ by $K^{\prime}$. Further, consider the extension $LK^{\prime}/K^{\prime}$ which has a base change of $K^{\prime}$ by $K^{\prime\prime}$ such that $K^{\prime\prime}/K$ is Galois. Then the extension $L/K$ has a base change of $K$ by $K^{\prime\prime}$.
Proof.
We have $K^{\prime}/K$ Galois with $\tilde{L}\cap K^{\prime}=K$ and $K^{\prime\prime}/K^{\prime}$ Galois with $\tilde{LK^{\prime}}\cap K^{\prime\prime}=K^{\prime}$. Observe that $\tilde{L}=\tilde{L}\cap\tilde{LK^{\prime}}$. Thus $\tilde{L}\cap K^{\prime\prime}=(\tilde{L}\cap\tilde{LK^{\prime}})\cap K^{\prime\prime}=\tilde{L}\cap(\tilde{LK^{\prime}}\cap K^{\prime\prime})=\tilde{L}\cap K^{\prime}=K$. Since $K^{\prime\prime}/K$ is Galois, we are done. ∎
8.1. A base change theorem for strong and weak cluster magnification
First we see a result about field extensions. The following lemma is a special case of [1, Prop. 2.5].
@@ 8.1. A base change theorem for strong and weak cluster magnification
Suppose $M/K$ is obtained by nontrivial strong cluster magnification from an $L/K$ with magnification factor $d$. Let $M/K$ have a base change of $K$ by $K^{\prime}$. Then $MK^{\prime}/K^{\prime}$ is obtained by nontrivial strong cluster magnification from $LK^{\prime}/K^{\prime}$ with the same magnification factor $d$.
@@
    Let $F^{\prime}:=FK^{\prime}$. Since, $F/K$ is Galois, $F^{\prime}/K^{\prime}$ is Galois. Also, $[F^{\prime}:K^{\prime}]=[F:K]=d>1$. Let $\tilde{L^{\prime}}$ be Galois closure of $L^{\prime}$ over $K^{\prime}$ in $\bar{K}$. Now, $N:=\tilde{L}K^{\prime}$ is Galois over $K^{\prime}$ and $L^{\prime}\subset N$. Hence, $\tilde{L^{\prime}}\subset N$. Since $F$ is linearly disjoint with $\tilde{L}$ over $K$, we have $\tilde{L}\cap F=K$. Also $\tilde{L}\cap K^{\prime}=K$ and $F\cap K^{\prime}=K$. Since $\tilde{L},F$ and $K^{\prime}$ are Galois over $K$. From Lem. 8.1.1 for $E_{1}=\tilde{L},E_{2}=F,E_{3}=K^{\prime}$, we have $\tilde{L}F\cap K^{\prime}=K\iff\tilde{L}K^{\prime}\cap FK^{\prime}=K^{\prime}$. So we have $N\cap F^{\prime}=K^{\prime}$. Thus, $\tilde{L^{\prime}}\cap F=K^{\prime}$. Hence $\tilde{L^{\prime}}$ and $F^{\prime}$ are linearly disjoint over $K^{\prime}$.
@@
Lem. 8.1.7 gives an alternate proof for Thm. 8.1.2 by using the criterion in Thm. 4.1.6.
We conclude the following result for the strong cluster magnification of polynomials from Thm. 8.1.2.
@@
Suppose $g$ over $K$ is obtained by nontrivial strong cluster magnification from an $f$ over $K$ with magnification factor $d$. Let $K_{g}/K$ have a base change of $K$ by $K^{\prime}$. Then $g$ over $K^{\prime}$ is obtained by nontrivial strong cluster magnification from $f$ over $K^{\prime}$ with the same magnification factor $d$.
@@
Suppose $M/K$ is obtained by weak cluster magnification from an $L/K$ with magnification factor $d$. Let $M/K$ have a base change of $K$ by $K^{\prime}$. Then $MK^{\prime}/K^{\prime}$ is obtained by weak cluster magnification from $LK^{\prime}/K^{\prime}$ with the same magnification factor $d$.
@@
Follows from last assertion in Lemma 8.1.7. ∎